In performing a chi-square goodness-of-fit test for a normal distribution, a researcher wants to make sure that all of the expected cell frequencies are at least five. The sample is divided into 7 intervals. The second through the sixth intervals all have expected cell frequencies of at least five. The first and the last intervals have expected cell frequencies of 1.5 each. After adjusting the number of intervals, the degrees of freedom for the chi-square statistic is ____.

2, 3,5, or 7
Answer #1

number of estimated variables =2 (mean and standard deviation)

here as 1st and last interval needs to be added to other intervals which makes number of categories =7-2 =5

therefore degrees of freedom for the chi-square statistic is =categories -number of estimated variables -1

=5-2-1 =2